# Mixed-State Measurement-Induced Phase Transitions in Imaginary-Time Dynamics

### v2

@@ Acknowledgements
We thank G.-Y. Zhu, B. Miao and Z. Bi for helpful discussions. Zenan Liu thanks the China Postdoctoral Science Foundation under Grants No.2024M762935 and NSFC Special Fund for Theoretical Physics under Grants No.12447119. Zhe Wang thanks the China Postdoctoral Science Foundation under Grants No.2024M752898. This project is supported by the Scientific Research Project (No.WU2024B027) and the Start-up Funding of Westlake University. The authors thank the high-performance computing center of Westlake University and the Beijing PARATERA Tech Co.,Ltd. for providing HPC resources.
Data availability. Data are available on Zenodo [zenodo].
@@ Appendix A A brief review of the stochastic series expansion quantum Monte Carlo
    $P_{\mathrm{remove}}(n+1\rightarrow n)=\min\bigg\{\frac{M-n+1}{\beta N_{b}\langle\alpha|H_{b}|\alpha\rangle},1\bigg\}$ (13)
The off-diagonal update will change the number of operators and update between diagonal operators and off-diagonal operators. To improve the update efficiency, some nonlocal update scheme such as the loop update, cluster update, or directed loop update are typically used [Sandvik2010lecture, sandvik2003stochastic, Anders2002Quantum].
Next, we take the TFIM in Sec. IV as an example to illustrate the SSE method, and it is similar to discuss the CDHM in Sec. V. The Hamiltonian of the TFIM is given by
$H=-J\sum_{\langle ij\rangle}Z_{i}Z_{j}-h\sum_{i}X_{i}$ (14)
@@
$H_{0,i}$ $\equiv hI_{i}$
$H_{1,i}$ $\equiv hX_{i}$
$H_{2,\langle ij\rangle}$ $\equiv J(Z_{i}Z_{j}+I_{i}I_{j})$ (15)
@@ Appendix B Quantum Monte Carlo algorithm for simulating the generalized partition function $\mathcal{Q}_{n_{d}}$
Figure 12: The measurement-averaged state $\overline{\rho}_{1}$ corresponds to the sampling configuration in the SSE stimulation, taking two circuit layers as an example. White rectangle and blue rectangle represents the diagonal operator and site operator respectively, while red rectangle denotes the auxiliary identity operators. In (a), the simulation performs a measurement operation after a time interval of $\tau/2$. At this moment, the manifold closes, and the two circuit layers are adhered together. In (b), no measurement is performed at $\tau/2$, which the manifold remains open, and the two circuits form a single larger layer. The dashed lines represent the periodic boundary conditions in imaginary time. The two subfigures correspond respectively to the two extreme cases of measurement rates $p=1$ and $p=0$.
As discussed in Sec. III.4, for a general measurement rate $p$, we only need to switch the simulated ensemble between Fig. 11(a) and (b) when the circuit layer is two. Specifically, before each Monte Carlo step, we consider the following merge-split process:
  1. (1)
    $P_{\mathrm{merge}}=\min\bigg\{\frac{p}{1-p},1\bigg\}$ (17)
  2. (2)
    $P_{\mathrm{split}}=\min\bigg\{\frac{1-p}{p},1\bigg\}$ (18)
The merge-split process can be naturally generalized to the case when the circuit layer is greater than two, i.e., more replicas of $e^{-\tau H}$, following the dicussions in Sec. III.5. For the Heisenberg model, the diagonal update, merge–split process, and nonlocal update are implemented in a manner analogous to those used for the TFIM under measurement. The cluster update scheme can be employed to perform updates between diagonal and off-diagonal operators in the Heisenberg model.
Appendix C Cluster formations and long-range orders
@@
In this section, we provide an intuitive, QMC-based picture for how local projective measurements enhance collective behavior in the system, ultimately generating a mixed ordered phase.
@@
As introduced in Appendix. A, the nonlocal cluster update is a crucial step to ensure the ergodicity of the QMC algorithm for the TFIM. Therefore, the size of the automatically formed clusters reflects the degree of collective behavior among spins: large clusters emerge when many spins fluctuate coherently. For example, in the 1D TFIM, when the transverse field is very strong $(h\to\infty)$, spins are dominated by the external field, and cluster updates produce only small, local clusters. Conversely, when the field is very weak, $(h\to 0)$, the Ising interactions dominate, connecting all sites into a single large clusters. In this regime, cluster construction naturally produces system-spanning clusters, reflecting the collective alignment of spins in the ferromagnetic phase.
@@
According to the merge-split process introduced in Appendix. B, the presence of measurements allows update lines to connect across replicas, effectively merging clusters from different replicas. Therefore, starting with the paramagnetic ground state of the TFIM, even local measurements can induce the formation of larger clusters during the QMC updates, which is similar to increasing the Ising coupling strength for the ground state of the TFIM. This is the reason why measurements can drive the system into a mixed-ordered phase with long-range order, from the perspective of cluster formations in QMC simulations. Similar reasoning also applies to the CDHM, where a cluster update is also used.
@@
Appendix D Results on the absence of off-diagonal long-range order in the 2D CDHM
[…]
Figure 13: Finite size extrapolation of off-diagonal spin-spin correlation $C_{XY}(L/2,L/2)$ at different measurement probability $p$ with $\tau=16$ (a) and $\tau=100$ (b).
In this section, we present additional numerical results demonstrating the absence of off-diagonal long-range order in the 2D CDHM under the MDITE protocol, as discussed in Sec. V. Combining the algorithm in Sec. C and the QMC tomography method [mao2025sampling, hari2025magic, chincholi2025detecting], we compute the correlation function

### v3

@@ Acknowledgements
We thank G.-Y. Zhu, B. Miao, Y. Guo, Z. Bi, C. Wang, and Y. Zou for helpful discussions. Zenan Liu thanks the China Postdoctoral Science Foundation under Grants No.2024M762935 and NSFC Special Fund for Theoretical Physics under Grants No.12447119. Zhe Wang thanks the China Postdoctoral Science Foundation under Grants No.2024M752898. This project is supported by the Scientific Research Project (No.WU2025B011) and the Start-up Funding of Westlake University. The authors thank the high-performance computing center of Westlake University for providing HPC resources.
Data availability. Data are available on Zenodo [21].
@@ Appendix A A brief review of the stochastic series expansion quantum Monte Carlo
    $P_{\mathrm{remove}}(n+1\rightarrow n)=\min\bigg\{\frac{M-n+1}{\beta N_{b}\langle\alpha|H_{b}|\alpha\rangle},1\bigg\}.$ (29)
The off-diagonal update will change the number of operators and update between diagonal operators and off-diagonal operators. To improve the update efficiency, some nonlocal update scheme such as the loop update, cluster update, or directed loop update are typically used [81, 80, 92].
Next, we take the TFIM in Sec. IV as an example to illustrate the SSE method, and it is similar to discuss the CDHM in Sec. VI. The Hamiltonian of the TFIM is given by
$H=-J\sum_{\langle ij\rangle}Z_{i}Z_{j}-h\sum_{i}X_{i}.$ (30)
@@
$H_{0,i}$ $\equiv hI_{i},$
$H_{1,i}$ $\equiv hX_{i},$
$H_{2,\langle ij\rangle}$ $\equiv J(Z_{i}Z_{j}+I_{i}I_{j}).$ (31)
@@ Appendix B Quantum Monte Carlo algorithm for simulating the generalized partition function $\mathcal{Q}_{n_{d}}$
Figure 15: The measurement-averaged state $\overline{\rho}_{1}$ corresponds to the sampling configuration in the SSE stimulation, taking two time steps as an example. White rectangle and blue rectangle represents the diagonal operator and site operator respectively, while red rectangle denotes the auxiliary identity operators. In (a), the simulation performs a measurement operation after a time interval of $\tau/2$. At this moment, the manifold closes, and the two time steps are adhered together. In (b), no measurement is performed at $\tau/2$, which the manifold remains open, and the two circuits form a single larger layer. The dashed lines represent the periodic boundary conditions in imaginary time. The two subfigures correspond respectively to the two extreme cases of measurement rates $p=1$ and $p=0$.
As discussed in Sec. III.4, for a general measurement rate $p$, we only need to switch the simulated ensemble between Fig. 14(a) and (b) when the number of time steps is two. Specifically, before each Monte Carlo step, we consider the following merge-split process:
  1. (1)
    $P_{\mathrm{merge}}=\min\bigg\{\frac{p}{1-p},1\bigg\};$ (33)
  2. (2)
    $P_{\mathrm{split}}=\min\bigg\{\frac{1-p}{p},1\bigg\}.$ (34)
The merge-split process can be naturally generalized to the case when the number of time steps is greater than two, i.e., more replicas of $e^{-\tau H}$, following the dicussions in Sec. III.5. For the Heisenberg model, the diagonal update, merge–split process, and nonlocal update are implemented in a manner analogous to those used for the TFIM under measurement. The cluster update scheme can be employed to perform updates between diagonal and off-diagonal operators in the Heisenberg model.
Appendix C Continuous-time limit of the 1D model
In this appendix, we perturbatively analyze the limit that $\tau\to 0$, $\tilde{\tau}\sim O(1)$, and $h\sim O(1)$. In this limit, Eq. (18) can be simplified using the BCH formula,
$e^{\tau A}e^{B}=e^{B+\tau A+\frac{\tau}{2}[A,B]+\frac{\tau}{12}[B,[B,A]]+O(\tau^{2})},$ (35)
where we identify
$A\equiv-\tilde{H},\qquad B\equiv\tilde{\tau}\sum_{i}Z_{i}^{L}Z_{i}^{R}.$ (36)
By dropping terms of order $O(\tau^{2})$ and higher, we have
$e^{\tau A}e^{B}\approx e^{-H_{\text{eff}}},$ (37)
with the effective Hamiltonian
$H_{\text{eff}}\equiv-B-\tau A-\frac{\tau}{2}[A,B]-\frac{\tau}{12}[B,[B,A]].$ (38)
For the two-copy TFIM Hamiltonian,
$\tilde{H}=-\frac{1}{2}\sum_{\langle ij\rangle}(Z_{i}^{L}Z_{j}^{L}+Z_{i}^{R}Z_{j}^{R})-\frac{h}{2}\sum_{i}(X_{i}^{L}+X_{i}^{R}),$ (39)
one finds, after straightforward algebra,
$A,B]=-\tilde{\tau}h\sum_{i}\bigg(iY_{i}^{L}Z_{i}^{R}+iZ_{i}^{L}Y_{i}^{R}\bigg),$ (40)
and
$B,[B,A]]=2\tilde{\tau}^{2}h\sum_{i}(X_{i}^{L}+X_{i}^{R}).$ (41)
Substituting these results into Eq. (38), we may rewrite the effective Hamiltonian, up to an overall rescaling by $\tilde{\tau}$, as
$H_{\text{eff}}=H_{Z}+(H_{ZZ}+H_{X}+H_{YZ}),$ (42)
where
$H_{Z}\equiv$ $-\sum_{i}Z_{i}^{L}Z_{i}^{R},$ (43)
$H_{ZZ}\equiv$ $-\alpha_{ZZ}\sum_{\langle ij\rangle}(Z_{i}^{L}Z_{j}^{L}+Z_{i}^{R}Z_{j}^{R}),$ (44)
$H_{X}\equiv$ $-\alpha_{X}\sum_{i}(X_{i}^{L}+X_{i}^{R}),$ (45)
$H_{YZ}\equiv$ $\alpha_{YZ}\sum_{i}(iY_{i}^{L}Z_{i}^{R}+iZ_{i}^{L}Y_{i}^{R}).$ (46)
The couplings are given by
$\alpha_{ZZ}\equiv\frac{\tau}{2\tilde{\tau}},\quad\alpha_{X}\equiv\frac{\tau}{\tilde{\tau}}\bigg(\frac{h}{2}+\frac{\tilde{\tau}^{2}h}{6}\bigg),\quad\alpha_{YZ}\equiv\frac{\tau h}{2}.$ (47)
Moreover, the effective Hamiltonian can be decomposed as
$H_{\mathrm{eff}}=H_{0}+H_{1},$ (48)
where $H_{0}\equiv H_{Z}$ and
$H_{1}\equiv\frac{\tau}{\tilde{\tau}}(H_{ZZ}+H_{X}+H_{YZ})$ (49)
is treated as a perturbation. We remark that, in this case, taking $\tau\to 0$ does not imply turning off the dynamics, but instead corresponds to a Trotterization of the combined evolution and measurement processes.
The unperturbed Hamiltonian $H_{0}$, whose ground-state energy is $E_{0}=-N$, dominates the low-energy Hilbert space and enforces the constraint $Z_{i}^{L}Z_{i}^{R}=+1$ for all $i$. Moreover, a single constraint violation at some site $j$, characterized by $Z_{j}^{L}Z_{j}^{R}=-1$, costs an energy gap $\Delta=2$. In the limit $\tau/\tilde{\tau}\to 0$, transitions out of the low-energy subspace induced by $H_{1}$ are perturbatively suppressed.
Next, we derive the low-energy Hamiltonian of $H_{\text{eff}}$. For convenience, we introduce the projectors
$P\equiv\prod_{i}P_{i},\qquad Q\equiv 1-P,$ (50)
where
$P_{i}\equiv\frac{1+Z_{i}^{L}Z_{i}^{R}}{2}$ (51)
projects onto the subspace with locally aligned left and right spins. Applying a Schrieffer-Wolff (SW) transformation, the low-energy effective Hamiltonian up to second order reads
$H_{\mathrm{eff}}\approx PH_{0}P+H^{(1)}_{\text{eff}}+H^{(2)}_{\text{eff}}$, (52)
where
$H^{(1)}_{\text{eff}}$= $PH_{1}P,$ (53)
$H^{(2)}_{\text{eff}}$= $-PH_{1}Q\frac{1}{H_{0}-E_{0}}QH_{1}P.$ (54)
Since $H_{0}$ imposes a hard constraint, the low-energy physics is governed by the remaining two correction terms.
The first-order term is given by
$H^{(1)}_{\mathrm{eff}}=H_{ZZ}=-\alpha_{ZZ}\sum_{\langle ij\rangle}(Z_{i}^{L}Z_{j}^{L}+Z_{i}^{R}Z_{j}^{R}),$ (55)
which describes FM Ising interactions within each copy.
At the second order, we have
$H^{(2)}_{\mathrm{eff}}=-(\alpha_{X}^{2}-\alpha_{YZ}^{2})\sum_{i}X_{i}^{L}X_{i}^{R},$ (56)
up to an unimportant constant energy shift.
Collecting the leading contributions, the effective low-energy Hamiltonian within the constrained subspace $Z_{i}^{L}Z_{i}^{R}=+1$ takes the form
$H^{\prime}_{\mathrm{eff}}=-\tilde{J}\sum_{\langle ij\rangle}(Z_{i}^{L}Z_{j}^{L}+Z_{i}^{R}Z_{j}^{R})-\tilde{h}\sum_{i}X_{i}^{L}X_{i}^{R},$ (57)
where the effective couplings are given by
$\tilde{J}\equiv\;\alpha_{ZZ}=\frac{\tau}{2\tilde{\tau}}$, (58)
and
$\tilde{h}\equiv\alpha_{X}^{2}-\alpha_{YZ}^{2}=\tau^{2}h^{2}\bigg[\frac{1}{\tilde{\tau}^{2}}\bigg(\frac{1}{2}+\frac{\tilde{\tau}^{2}}{6}\bigg)^{2}-\frac{1}{4}\bigg]>0.$ (59)
Inside the constrained subspace $Z_{i}^{L}Z_{i}^{R}=+1$, the constrained Hilbert space is exactly that of a single qubit per site. Therefore, by defining effective Pauli operators
$\tilde{Z}_{i}\equiv\tilde{Z}_{i}^{L}\equiv\tilde{Z}_{i}^{R},\quad\tilde{X}_{i}\equiv X_{i}^{L}X_{i}^{R},$ (60)
which satisfy the standard Pauli algebra, we have
$H^{\prime}_{\mathrm{eff}}=-2\tilde{J}\sum_{\langle ij\rangle}\tilde{Z}_{i}\tilde{Z}_{j}-\tilde{h}\sum_{i}\tilde{X}_{i}.$ (61)
This Hamiltonian is precisely a single-copy TFIM. For the 1D case, it has a quantum critical point at $2\tilde{J}=\tilde{h}$ according to the Kramers-Wannier duality, which belongs to the 2D classical Ising universality class.
Furthermore,
$2\tilde{J}\gtrless\tilde{h}\quad\Longleftrightarrow\quad\frac{\tau}{\tilde{\tau}}$ $\gtrless\tau^{2}h^{2}\left[\frac{1}{\tilde{\tau}^{2}}\left(\frac{1}{2}+\frac{\tilde{\tau}^{2}}{6}\right)^{2}-\frac{1}{4}\right]$ (62)
@@ Appendix C Continuous-time limit of the 1D model
Since $\tau\to 0$, the left-hand side of Eq. (62) becomes parametrically large. Consequently, by increasing $\tilde{\tau}$, or equivalently the measurement rate $p$, the system can be driven across the critical point, inducing a quantum phase transition from the FM ordered phase to the PM disordered phase.
This behavior is physically intuitive. In the continuous time limit, when the measurement rate is much larger than the ITE scale, i.e., $\tilde{\tau}\gg\tau$, the steady state is dominated by the decoherence generated by $e^{\tilde{\tau}\sum_{i}Z_{i}^{L}Z_{i}^{R}}$ in Eq. (18). In this regime, the measurements strongly constrain the system, preventing the ITE from establishing long-range order, even $h<1$.
The same conclusion follows directly from the effective Hamiltonian in Eq. (42). In the extreme limit $\tau=0$, the dynamics is entirely governed by the measurement process, and the steady state coincides with the ground state of $H_{\mathrm{eff}}=-\sum_{i}Z_{i}^{L}Z_{i}^{R}$. This state enforces local alignment between the two copies at each site but does not generate spatial Ising order within either copy.
We emphasize that this mechanism is distinct from that observed in Sec. IV away from the limit, where increasing the measurement rate $p$ instead drives a phase transition from a disordered phase to an ordered phase.
Appendix D Weak-measurement limit of the 1D model
In the weak-measurement limit, $\tilde{\tau}\to 0\sim p/2$, with $\tau\sim O(1)$ and $h\sim O(1)$, one can similarly derive an effective two-copy Hamiltonian
$H_{\text{eff}}=H_{0}+H_{1},$ (63)
where
$H_{0}\equiv\tilde{H}=-\frac{1}{2}\sum_{\langle ij\rangle}(Z_{i}^{L}Z_{j}^{L}+Z_{i}^{R}Z_{j}^{R})-\frac{h}{2}\sum_{i}(X_{i}^{L}+X_{i}^{R}),$ (64)
and
$H_{1}\equiv H_{Z}+H_{YZ}+H_{XZZ}+H_{Y}$ (65)
with
$H_{Z}\equiv$ $\bigg(\frac{p\tau h}{12}-\frac{p}{2\tau}\bigg)\sum_{i}Z_{i}^{L}Z_{i}^{R},$ (66)
$H_{YZ}\equiv$ $-\frac{ph}{4}\sum_{i}\bigg(iY_{i}^{L}Z_{i}^{R}+iZ_{i}^{L}Y_{i}^{R}\bigg),$ (67)
$H_{XZZ}\equiv$ $-\frac{p\tau h}{24}\sum_{\langle ij\rangle}(X_{i}^{L}Z_{j}^{L}Z_{i}^{R}$
$+Z_{i}^{L}X_{j}^{L}Z_{j}^{R}+Z_{i}^{L}X_{i}^{R}Z_{j}^{R}+Z_{j}^{L}Z_{i}^{R}X_{j}^{R},)$ (68)
$H_{Y}\equiv$ $-\frac{p\tau h}{12}\sum_{i}Y_{i}^{L}Y_{i}^{R}.$ (69)
The unperturbed Hamiltonian $H_{0}$ has a global $\mathbb{Z}_{2}^{L}\times\mathbb{Z}_{2}^{R}$ symmetry. Accordingly, the projector onto the ground-state sector can be written as
$P=\prod_{i}P_{i},\quad P_{i}\equiv\frac{1+X_{i}^{L}}{2}\frac{1+X_{i}^{R}}{2}.$ (70)
[…]
This perturbative stability originates not only from the finite excitation gap, but also from the product structure of the ground-state projector $P=\prod_{i}P_{i}$, which imposes simultaneous constraints on the two copies. As a result, any local perturbation necessarily creates gapped excitations and cannot generate nontrivial operators within the low-energy subspace at these orders.
Therefore, the weak-measurement limit also does not exhibit physics that is qualitatively consistent with the numerical results presented in Sec. IV.
Appendix E Cluster formations and long-range orders
@@
In this appendix, we provide an intuitive, QMC-based picture for how local projective measurements enhance collective behavior in the system, ultimately generating a mixed ordered phase.
@@
As introduced in Appendix. A, the nonlocal cluster update is a crucial step to ensure the ergodicity of the QMC algorithm for the TFIM. Therefore, the size of the automatically formed clusters reflects the degree of collective behavior among spins: large clusters emerge when many spins fluctuate coherently. For example, in the 1D TFIM, when the transverse field is very strong $(h\to\infty)$, spins are dominated by the external field, and cluster updates produce only small, local clusters. Conversely, when the field is very weak, $(h\to 0)$, the Ising interactions dominate, connecting all sites into a single large clusters. In this regime, cluster construction naturally produces system-spanning clusters, reflecting the collective alignment of spins in the FM phase.
@@
According to the merge-split process introduced in Appendix. B, the presence of measurements allows update lines to connect across replicas, effectively merging clusters from different replicas. Therefore, starting with the PM ground state of the TFIM, even local measurements can induce the formation of larger clusters during the QMC updates, which is similar to increasing the Ising coupling strength for the ground state of the TFIM. This reflects why measurements can drive the system into a mixed-ordered phase with long-range order, from the perspective of cluster formations in QMC simulations. Similar reasoning also applies to the CDHM, where a cluster update is also used.
@@
Appendix F Fittings of the other two dynamical exponents for the 1D TFIM
In this appendix we present additional numerical results for the extraction of the dynamical exponent $z$ in the example of 1D TFIM in Sec. IV.5, using Eq. (12). These results complement the analysis presented in Sec. IV.5 and further support the conclusion that the mixed-state critical points exhibit a dynamical exponent close to $z=1$. See Table 4.
$(\tau,h,p)$ $L_{\min}$ $a$ $z$
(1.2,1.8373,0.8) 64 9.2(2) 1.013(5)
72 9.1(3) 1.010(6)
80 9.0(3) 1.009(7)
88 8.6(4) 0.997(9)
96 8.1(5) 0.99(1)
(0.2651,2.5,0.5) 64 2.2(1) 0.97(1)
72 2.2(1) 0.97(1)
80 2.0(1) 0.96(1)
88 1.8(2) 0.93(2)
96 2.0(2) 0.96(3)
Table 4: Extraction of the dynamical exponent $z$ from the relaxation time for two critical points in the example of TFIM. System sizes range from L=64,72,,128L=64,72,...,128. To test the stability of the extracted exponent, small system sizes are gradually discarded. $L_{\min}$ denotes the smallest system size included in the fits.
Appendix G Dynamical scaling of the example of 2D CDHM
Figure 16: When setting $(\tau,g,p)=(1,3.5,0.3538)$ for the MDITE with the 2D CDHM: (a) The relaxation process of $\langle|m|\rangle$ with increasing the number of time steps $n_{d}$ for various system sizes $L$. It satisfies the formula $m=a\exp(-n_{d}/\xi_{\tau})+b$, where $a$, $\xi_{\tau}$ and $b$ are fitting parameters. (b) Extracting the dynamical exponent $z$ from $\xi_{\tau}\sim L^{z}$ with various system size $L$.
Similar to the example of 1D TFIM, we extract the dynamical exponent $z$ for the 2D CDHM by simulating the relaxation dynamics of the order parameter $\langle|m_{z}|\rangle$.
As a representative case, for $(\tau,g)=(1,3.5)$ the phase transition occurs at $p_{c}\approx 0.3538$, as shown in Fig. 12(a) and (b). From the relaxation behavior illustrated in Fig. 16, we extract the relaxation time $\xi_{\tau}$ and obtain a dynamical exponent $z\approx 0.963$, which is close to unity. The stability analysis further confirms $z\approx 1$, as summarized in Table 5.
We also determine the dynamical exponent for two additional parameter sets, $(\tau,g,p)=(2,2.806,0.3)$ and $(0.4679,3.0,0.1)$. The extracted values again yield $z\approx 1$.
$(\tau,g,p)$ $L_{\min}$ $a$ $z$
(1,3.5,0.3538) 12 4.63(6) 0.963(4)
16 4.56(7) 0.959(4)
20 4.64(8) 0.963(5)
24 4.6(1) 0.963(6)
(2,2.806,0.3) 8 11.3(3) 1.092(6)
12 11.3(3) 1.093(6)
16 11.2(3) 1.091(7)
20 11.4(3) 1.096(8)
(0.4679,3.0,0.1) 12 2.17(6) 1.015(8)
16 2.17(6) 1.015(8)
20 2.14(7) 1.011(8)
24 2.08(8) 1.00(1)
Table 5: Extraction of the dynamical exponent $z$ from the relaxation time in the CDHM using system sizes $L\in[8,48]$. Different parameter sets $(\tau,g,p)$ correspond to three quantum critical points. To test the stability of the fits, small system sizes are gradually discarded. The resulting values are consistently close to $z\approx 1$.
Appendix H Results on the absence of off-diagonal long-range order in the 2D CDHM
Figure 17: Finite size extrapolation of off-diagonal spin-spin correlation $C_{XY}(L/2,L/2)$ at different measurement probability $p$ with (a) $\tau=16$ and (b) $\tau=100$ .
In this appendix, we present additional numerical results demonstrating the absence of off-diagonal long-range order in the 2D CDHM under the MDITE protocol, as discussed in Sec. VI. Combining the algorithm in Sec. E and the QMC tomography method [57, 99, 12], we compute the correlation function